# Pistis-RAG: A Scalable Cascading Framework Towards Trustworthy Retrieval-Augmented Generation

### v3

@@ 4.1. RAG "Re-Rankers" in the Pre-Ranking Stage
While terms like $bge-reranker$ are commonly used for components within RAG systems, they can be misleading. Information Retrieval systems like search engines employ re-ranking models in a later stage to refine results based on specific business needs.
In RAG, however, these models are used upfront during the pre-ranking stage to narrow down candidate documents. Therefore, a more accurate term for this function would be "pre-ranker."
@@ 5.2. Datasets
In assessing the effectiveness of the Pistis-RAG framework, it is essential to leverage diverse datasets that offer a comprehensive evaluation of its capabilities. MMLU is an invaluable resource, providing insights into the framework’s adaptability and performance across various domains.
The MMLU dataset is a valuable resource for researchers developing speech recognition and natural language processing systems. The data set has been used to train various speech recognition systems and has been shown to improve the performance of these systems. The dataset has also been used to develop natural language processing systems that can understand and generate spoken language.
With a total of 15,908 questions, MMLU provides a robust benchmark for evaluating language models’ performance across various tasks. Particularly noteworthy is its relevance to healthcare, making it an ideal choice for assessing the Pistis-RAG framework’s ability to handle complex medical queries.
It evaluates a framework’s ability to understand and generate coherent responses across various subjects, making it an ideal choice for evaluating the Pistis-RAG framework’s performance in handling complex queries.
@@ 5.3. Experimental Setup
Our experimental setup is meticulously designed to optimize the performance of our models in both retrieval and response generation. Below are the key components of our setup:
To evaluate the effectiveness of our proposed method, we conducted experiments on the MMLU dataset. We first built an index of the MMLU training set, then used BEG-M3 to retrieve the top 10 candidate few-shots based on Milvus. Then, we used BEG-reranker-larger to pre-rank the candidate few-shots and obtain the top five candidate few-shots. Finally, we used these five candidate few-shots to generate 5-shot results as the baseline.
Matching: We used BEG-M3 to retrieve the top 10 candidate few-shots based on Milvus. BEG-M3 is a dense vector similarity search engine that can efficiently retrieve similar vectors from a large collection of vectors. Milvus is a vector database that stores and manages vectors.
@@ 6. Conclusions
This study highlights cascade modeling and optimization as critical areas for robust large-scale online AI-generated content (AIGC) systems.
We revisited the Retrieval-Augmented Generation problem through a content-centric lens, uncovering potential shortcomings in large-scale deployments. Firstly, a mismatch exists between the intended and actual function of the "re-ranker" component, suggesting it acts more like a pre-ranker. We addressed this by proposing distinct ranking and re-ranking stages within RAG and a listwide ranking approach. Furthermore, we introduced reasoning and aggregation stages specific to AIGC services, paving the way for a more comprehensive and high-performing system.

### v4

@@ 4.1. RAG "Re-Rankers" in the Pre-Ranking Stage
While terms like $bge-reranker$ are commonly used for components within RAG systems, they can be misleading. Information Retrieval systems like search engines employ re-ranking models in a later stage to refine results based on specific business needs. In RAG, however, these models are used upfront during the pre-ranking stage to narrow down candidate documents. Therefore, a more accurate term for this function would be "pre-ranker."
@@ 5.2. Datasets
In assessing the effectiveness of the Pistis-RAG framework, it is essential to leverage datasets that offer a comprehensive evaluation of its capabilities. MMLU is an invaluable resource, providing insights into the framework’s adaptability and performance across various domains. With 15,908 questions, MMLU provides a robust benchmark for evaluating language models’ performance across multiple tasks. Particularly noteworthy is its relevance to healthcare, making it an ideal choice for assessing the Pistis-RAG framework’s ability to handle complex queries. It evaluates a framework’s ability to understand and generate coherent responses across various subjects, making it an ideal choice for evaluating the Pistis-RAG framework’s performance in handling complex queries.
@@ 5.3. Experimental Setup
Our experimental setup is meticulously designed to optimize the performance of our models in both retrieval and response generation. Below are the key components of our setup: To evaluate the effectiveness of our proposed method, we conducted experiments on the MMLU dataset. We first built an index of the MMLU training set, then used BEG-M3 to retrieve the top 10 candidate few-shots based on Milvus. Then, we used BEG-reranker-larger to pre-rank the candidate few-shots and obtain the top five candidate few-shots. Finally, we used these five candidate few-shots to generate 5-shot results as the baseline.
Matching: We used BEG-M3 to retrieve the top 10 candidate few-shots based on Milvus. BEG-M3 is a dense vector similarity search engine that can efficiently retrieve similar vectors from many vectors. Milvus is a vector database that stores and manages vectors.
@@ 6. Conclusions
This study highlights cascade modeling and optimization as critical areas for robust large-scale online AIGC systems. We revisited the Retrieval-Augmented Generation problem through a content-centric lens, uncovering potential shortcomings in large-scale deployments. Firstly, a mismatch exists between the intended and actual function of the "re-ranker" component, suggesting it acts more like a pre-ranker. We addressed this by proposing distinct ranking and re-ranking stages within RAG and a listwide ranking approach. Furthermore, we introduced reasoning and aggregation stages specific to AIGC services, paving the way for a more comprehensive and high-performing system.